OpenStudy (anonymous):

A 1600 principal earns 7% interest, compounded semi-annually twice per year. After 33 years, what is the balance in the account?

OpenStudy (retireed):

Do you have a formula?

OpenStudy (anonymous):

No

OpenStudy (retireed):

Why not? Surely your teacher gave you some sort of lesson on this subject or at least gave you book or study notes.

OpenStudy (anonymous):

Maybe, I've looked over them but I'm not really understanding the whole thing

OpenStudy (retireed):

Okay fair enough. A is the amount you want to find P is the principal P = 1600 Semi-annually means 2 interest periods per year n =2 t is the number of years t = 33 N is the number of interest periods N = t * n N = 33 * 2 N = 66 I is the interest percentage I = 7% i is the interest rate for the interest period i = I/100/n = 7%/100/2 i =0.035

OpenStudy (retireed):

Now I need to check the formula to make sure I have the right one. I think it is A = P ( 1+ i)^N Does this look familiar?

OpenStudy (anonymous):

Yes thank you